OpenStudy (anonymous):

Given f(x)=√x and g(x)=x^4, find g o f and its domain

OpenStudy (apoorvk):

g o f = g(f(x)) so here, g o f = sqrt(x^4) =x^2 domain of GoF--> all R.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

domain of \(f\) is \(x\geq 0\) and so is the domain of \(g\circ f\)

OpenStudy (anonymous):

while it is true that \[g\circ f(x)=x^2\] the domain is the domain of \(f\)

OpenStudy (apoorvk):

@satellite73 so it means that even though g 0 f is a totally new function, the constraints from the constituent functions shall apply. right?

OpenStudy (anonymous):

yes, if you cannot evaluate \(f\) you cannot evaluate \(g\circ f\)

OpenStudy (apoorvk):

okay, thanks. and sorry @rose1285 for the wrong info.

OpenStudy (anonymous):

since \(g\circ f\) means first do \(f\) then do \(g\)

OpenStudy (anonymous):

Thanks so much guys!